### v1

@@ Appendix C Finite Difference Monte Carlo Estimators
The Finite Difference Monte Carlo strategy applied to the prediction risk estimate $\mathsf{APURE}^{\mathcal{P}}$ of Equation (LABEL:eq:apure) consists in rewriting the second term in the definition of $\mathsf{APURE}^{\mathcal{P}}$ in Equation (LABEL:eq:apure) in a tractable way. The proof thus focuses on this term and aims at demonstrating that
$\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\sum_{t=1}^{T}\widehat{\mathsf{X}}_{t}^{-t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\Psi_{t}(\boldsymbol{\mathsf{Y}})\mathsf{Y}_{t}\right]=\mathbb{E}_{\boldsymbol{\mathsf{Y}},\boldsymbol{\zeta}}\left[\sum_{t=1}^{T}\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\Psi_{t}(\boldsymbol{\mathsf{Y}})\mathsf{Y}_{t}\right.$
$\quad-\left\langle\mathrm{diag}(\boldsymbol{\alpha}\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}}))\partial_{\boldsymbol{\mathsf{Y}}}\widehat{\boldsymbol{\mathsf{X}}}[\boldsymbol{\zeta}],\mathrm{diag}(\boldsymbol{\mathsf{Y}})\boldsymbol{\zeta}\right\rangle\left],{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}\sum_{t=1}^{T}}\right.$ (74)
@@
$\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\widehat{\mathsf{X}}_{t}^{-t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\Psi_{t}(\boldsymbol{\mathsf{Y}})\mathsf{Y}_{t}\right]=$ (75)
$\mathbb{E}_{\boldsymbol{\mathsf{Y}},\boldsymbol{\zeta}}\left[\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\Psi_{t}(\boldsymbol{\mathsf{Y}})\mathsf{Y}_{t}-\alpha_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\left(\partial_{\boldsymbol{\mathsf{Y}}}\widehat{\boldsymbol{\mathsf{X}}}[\boldsymbol{\zeta}]\right)_{t}\mathsf{Y}_{t}\zeta_{t}\right]$
@@
$\begin{split}&\mathbb{E}_{\boldsymbol{\zeta}}\left[\widehat{\mathsf{X}}_{t}^{-t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\Psi_{t}(\boldsymbol{\mathsf{Y}})\mathsf{Y}_{t}\right]\\ \underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}&\mathbb{E}_{\boldsymbol{\zeta}}\left[\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}})\Psi_{t}(\boldsymbol{\mathsf{Y}})\mathsf{Y}_{t}-\alpha_{t}\left(\partial_{\boldsymbol{\mathsf{Y}}}\widehat{\boldsymbol{\mathsf{X}}}[\boldsymbol{\zeta}]\right)_{t}\zeta_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\mathsf{Y}_{t}\right].\end{split}$ (78)
@@
Assuming that $\forall t\in\{1,\ldots,T\},\,\forall\boldsymbol{\mathsf{Y}}\in\mathbb{R}^{T},\,\Psi_{t}(\boldsymbol{\mathsf{Y}})\neq 0$, a similar proof holds for the estimation risk estimate. Applying the Finite Difference Monte Carlo strategy applied to the estimation risk estimate $\mathsf{APURE}^{\mathcal{E}}$ of Equation (35) amounts to rewrite the second term in the definition of $\mathsf{APURE}^{\mathcal{E}}$ in Equation (35) in a tractable way. The proof thus focuses on this term and aims at demonstrating that
$\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\sum_{t=1}^{T}\frac{\widehat{\mathsf{X}}_{t}^{-t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\mathsf{Y}_{t}\right]=\mathbb{E}_{\boldsymbol{\mathsf{Y}},\boldsymbol{\zeta}}\left[\sum_{t=1}^{T}\sum_{t=1}^{T}\frac{\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\mathsf{Y}_{t}\right.$
$\quad-\left\langle\mathrm{diag}(\boldsymbol{\alpha}\centerdot/\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}}))\,\partial_{\boldsymbol{\mathsf{Y}}}\widehat{\boldsymbol{\mathsf{X}}}[\boldsymbol{\zeta}],\mathrm{diag}(\boldsymbol{\mathsf{Y}})\boldsymbol{\zeta}\right\rangle\left],{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}\sum_{t=1}^{T}}\right.$ (79)
@@
$\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\frac{\widehat{\mathsf{X}}_{t}^{-t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\mathsf{Y}_{t}\right]=$ (80)
$\mathbb{E}_{\boldsymbol{\mathsf{Y}},\boldsymbol{\zeta}}\left[\frac{\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\mathsf{Y}_{t}-\frac{\alpha_{t}}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\left(\partial_{\boldsymbol{\mathsf{Y}}}\widehat{\boldsymbol{\mathsf{X}}}[\boldsymbol{\zeta}]\right)_{t}\mathsf{Y}_{t}\zeta_{t}\right].$

### v3

@@ Appendix C Finite Difference Monte Carlo Estimators
The Finite Difference Monte Carlo strategy applied to the prediction risk estimate $\mathsf{APURE}^{\mathcal{P}}$ of Equation (31) consists in rewriting the second term in the definition of $\mathsf{APURE}^{\mathcal{P}}$ in Equation (31) in a tractable way. The proof thus focuses on this term and aims at demonstrating that
$\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\sum_{t=1}^{T}\widehat{\mathsf{X}}_{t}^{-t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\Psi_{t}(\boldsymbol{\mathsf{Y}})\mathsf{Y}_{t}\right]=\mathbb{E}_{\boldsymbol{\mathsf{Y}},\boldsymbol{\zeta}}\left[\sum_{t=1}^{T}\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\Psi_{t}(\boldsymbol{\mathsf{Y}})\mathsf{Y}_{t}\right.\quad-\left\langle\mathrm{diag}(\boldsymbol{\alpha}\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}}))\partial_{\boldsymbol{\mathsf{Y}}}\widehat{\boldsymbol{\mathsf{X}}}[\boldsymbol{\zeta}],\mathrm{diag}(\boldsymbol{\mathsf{Y}})\boldsymbol{\zeta}\right\rangle\left],{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}\sum_{t=1}^{T}}\right.$ (69)
@@
$\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\widehat{\mathsf{X}}_{t}^{-t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\Psi_{t}(\boldsymbol{\mathsf{Y}})\mathsf{Y}_{t}\right]=\mathbb{E}_{\boldsymbol{\mathsf{Y}},\boldsymbol{\zeta}}\left[\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\Psi_{t}(\boldsymbol{\mathsf{Y}})\mathsf{Y}_{t}-\alpha_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\left(\partial_{\boldsymbol{\mathsf{Y}}}\widehat{\boldsymbol{\mathsf{X}}}[\boldsymbol{\zeta}]\right)_{t}\mathsf{Y}_{t}\zeta_{t}\right]$ (70)
@@
$\begin{split}\mathbb{E}_{\boldsymbol{\zeta}}\left[\widehat{\mathsf{X}}_{t}^{-t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\Psi_{t}(\boldsymbol{\mathsf{Y}})\mathsf{Y}_{t}\right]\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathbb{E}_{\boldsymbol{\zeta}}\left[\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}})\Psi_{t}(\boldsymbol{\mathsf{Y}})\mathsf{Y}_{t}-\alpha_{t}\left(\partial_{\boldsymbol{\mathsf{Y}}}\widehat{\boldsymbol{\mathsf{X}}}[\boldsymbol{\zeta}]\right)_{t}\zeta_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\mathsf{Y}_{t}\right].\end{split}$ (73)
@@
Assuming that $\forall t\in\{1,\ldots,T\},\,\forall\boldsymbol{\mathsf{Y}}\in\mathbb{R}^{T},\,\Psi_{t}(\boldsymbol{\mathsf{Y}})\neq 0$, a similar proof holds for the estimation risk estimate. Applying the Finite Difference Monte Carlo strategy to the estimation risk estimate $\mathsf{APURE}^{\mathcal{E}}$ of Equation (33) amounts to rewrite the second term in the definition of $\mathsf{APURE}^{\mathcal{E}}$ in Equation (33) in a tractable way. The proof thus focuses on this term and aims at demonstrating that
$\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\sum_{t=1}^{T}\frac{\widehat{\mathsf{X}}_{t}^{-t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\mathsf{Y}_{t}\right]=\mathbb{E}_{\boldsymbol{\mathsf{Y}},\boldsymbol{\zeta}}\left[\sum_{t=1}^{T}\sum_{t=1}^{T}\frac{\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\mathsf{Y}_{t}\right.\quad-\left\langle\mathrm{diag}(\boldsymbol{\alpha}\centerdot/\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}}))\,\partial_{\boldsymbol{\mathsf{Y}}}\widehat{\boldsymbol{\mathsf{X}}}[\boldsymbol{\zeta}],\mathrm{diag}(\boldsymbol{\mathsf{Y}})\boldsymbol{\zeta}\right\rangle\left],{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}\sum_{t=1}^{T}}\right.$
@@
$\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\frac{\widehat{\mathsf{X}}_{t}^{-t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\mathsf{Y}_{t}\right]=\mathbb{E}_{\boldsymbol{\mathsf{Y}},\boldsymbol{\zeta}}\left[\frac{\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\mathsf{Y}_{t}-\frac{\alpha_{t}}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\left(\partial_{\boldsymbol{\mathsf{Y}}}\widehat{\boldsymbol{\mathsf{X}}}[\boldsymbol{\zeta}]\right)_{t}\mathsf{Y}_{t}\zeta_{t}\right].$ (74)